OpenStudy (1969tabby):

Solve w=Cr^-2 for r. I know the r becomes 1/r^2, but I con not figure out how to get the rest of the equation.

OpenStudy (anonymous):

r=sqrt(c/w)

OpenStudy (anonymous):

:) \[r^{2}=C/w.....r=\sqrt{\frac{c}{w}}\]

OpenStudy (anonymous):

s\[\S ^{-1}\rightarrow 1/\S so R ^{-2} \rightarrow 1/R ^{2}\]

OpenStudy (anonymous):

C not included i think \[w = r ^{-2}\]

OpenStudy (1969tabby):

Thank you! I was really turned aroun on this one...